OpenStudy (anonymous):

Please Help Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5)______

OpenStudy (mathmale):

Altho' there may be more sophisticated ways of solving this problem, I'd just write out the first five terms of this sequence: f(1)=2 f(2)=3 f(3)=f(1)+f(2) + f(3-1) = 2 + 3 + 3 = 8 f(4)=f(1)+f(2) + f(4-1)=? and so on.

OpenStudy (anonymous):

Thank you so much for help @mathmale!

OpenStudy (anonymous):

is it 18? I found 18